OpenStudy (anonymous):

square root(x+27) +3=x

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Hello, my name is Stephen. I'll be your personal assistant. For further understanding, is the equation below the same equation you posted above? \[\sqrt{x + 27} + 3 = x\]

OpenStudy (anonymous):

yes

OpenStudy (compassionate):

Here at OpenStudy we guide our users to the answer. Do you know what the first step is?

OpenStudy (anonymous):

subtract 3

OpenStudy (compassionate):

Correct. \[\sqrt{x + 27} = x - 3\] What do you think we do now?

OpenStudy (anonymous):

I'm not trying to learn I just want the answer... to be frank

OpenStudy (compassionate):

x = 9 Please come back next time. If you have any further questions feel free to tag me.

OpenStudy (anonymous):

Can you factor a trinomial for me?

OpenStudy (compassionate):

Yes.

OpenStudy (anonymous):

3x^3y+23x^2y^2+14xy^3

OpenStudy (compassionate):

\[3x^3 y + 23x^2 y^2 + 14xy^3\] Is that it? ^

OpenStudy (anonymous):

yes

OpenStudy (compassionate):

xy(x + 7y)(3x + 2y)

OpenStudy (anonymous):

thank youuuu

OpenStudy (compassionate):

No problem. Come back enxt time.